two buses at rest, left at 4pm from two stations that are 5km apart and travelled towards each other .if the first bus constantly accelerates at 1m.s2 while the second bus constantly accelerates at 1.5m.s2, calculate the distance (in km) that the first bus must travel before the two buses pass each other


1
Expert's answer
2021-02-24T12:45:30-0500

"s_1=a_1t^2/2\to t^2=2s_1/a_1"


"s_2=a_2t^2/2\to 5000-s_1=a_2t^2/2 \to t^2=2(5000-s_1)/a_2"


"2s_1/a_1=2(5000-s_1)/a_2\to 2s_1/1=2(5000-s_1)/1.5\to"


"s_1=2000\ (m)" . Answer